Compute​ P(X) using the binomial probability formula. Then determine whether the normal distribution can be used to estimate this probability. If​ so, approximate​ P(X) using the normal distribution and compare the result with the exact probability.

n=56​, p=0.6​, and X=22
Answer #1

Solution

Given that ,

p = 0.6

1 - p = 1 - 0.6 = 0.4

n = 56

x = 22

Using binomial probability formula ,

P(X = x) = ((n! / x! (n - x)!) * px * (1 - p)n - x

P(X = 22) = ((56! / 22! (56 - 22)!) * 0.622 * (0.4)56 - 22

=  ((56! / 22! (34)!) * 0.622 * (0.4)34

= 0.0008

Probability = 0.0008